40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Entry  Fri Nov 15 17:29:14 2019, Raymond, Lab Infrastructure, General, Moved IR Labs cryostat into QIL from Cryo Lab IMG_0270.JPGIMG_0269.JPG
    Reply  Tue Nov 19 20:16:40 2019, Raymond, Lab Infrastructure, General, Moved IR Labs cryostat into QIL from Cryo Lab IMG_0483.JPG
       Reply  Wed Nov 20 11:59:17 2019, Koji, Lab Infrastructure, General, Moved IR Labs cryostat into QIL from Cryo Lab 
          Reply  Wed Nov 20 12:25:58 2019, Koji, Lab Infrastructure, General, Moved IR Labs cryostat into QIL from Cryo Lab 
             Reply  Wed Nov 20 17:44:10 2019, Koji, Lab Infrastructure, General, Moved IR Labs cryostat into QIL from Cryo Lab P_20191120_174034_vHDR_On.jpgP_20191120_174040_vHDR_On.jpg
                Reply  Thu Nov 21 21:45:10 2019, Raymond, Lab Infrastructure, General, Moved IR Labs cryostat into QIL from Cryo Lab DT02H_18_32PN.pdf
                   Reply  Fri Nov 22 13:21:26 2019, Raymond, Lab Infrastructure, General, Moved IR Labs cryostat into QIL from Cryo Lab IMG_0497.JPGIMG_0498.JPG
                      Reply  Fri Nov 22 15:33:40 2019, Koji, Lab Infrastructure, General, Moved IR Labs cryostat into QIL from Cryo Lab P_20191122_150845_vHDR_On.jpg
                         Reply  Sat Nov 23 13:00:06 2019, Koji, Lab Infrastructure, Optics, Optical window transmission measurement optical_window_T.pdfP_20191122_183426_vHDR_On.jpgP_20191122_183436_vHDR_On.jpg
                         Reply  Tue Nov 26 15:36:59 2019, Raymond, Lab Infrastructure, General, Moved IR Labs cryostat into QIL from Cryo Lab IRLabsQIL.pdf
                            Reply  Wed Dec 4 13:54:04 2019, Raymond, Lab Infrastructure, General, Moved IR Labs cryostat into QIL from Cryo Lab 
Message ID: 2461     Entry time: Tue Nov 26 15:36:59 2019     In reply to: 2457     Reply to this: 2467
Author: Raymond 
Type: Lab Infrastructure 
Category: General 
Subject: Moved IR Labs cryostat into QIL from Cryo Lab 

The IR Labs cryostat is now pumping down. The thor labs posts are necessary for centering the optical port at a height of 5.5". The orange cabinet used to elevate the pumping station was relocated from next to the computer (was, and still is, empty). The power supply is for the wide range vacuum gauge attached to the pumping line. 

I've ordered a liquid nitrogen dewar for arrival Monday morning.

For future reference, the gauge communicates via a VGA port, for which the pin-to-wire association is as such: 1-Black; 2-Brown; 3-Red; 4-Orange; 5-Yellow; 6-Green; 7-Blue; 8-Purple; 9-Grey; 10-White; 11-Pink; 12-Cyan; 13-Black/White; 14-Brown/White; 15-Red/White  (MKS 901P non-ethercat wiring diagram)

Attachment 1: IRLabsQIL.pdf  21.330 MB  | Hide | Hide all
IRLabsQIL.pdf
ELOG V3.1.3-